A bracelet that is suppose to be 18k gold weighs 238grams. Gold weighs 19.3 grams per cubic centimeter. Copper weighs 9 grams per cubic centimeter. The volume of the bracelet is 15 cubic centimeters. Is the bracelet really 18k? I know that isn't, I need to set it up in substitution method.
Answer by ankor@dixie-net.com(22740)

A bracelet that is suppose to be 18k gold weighs 238grams.
Gold weighs 19.3 grams per cubic centimeter.
Copper weighs 9 grams per cubic centimeter.
The volume of the bracelet is 15 cubic centimeters.
Is the bracelet really 18k?
:
pure gold is 24 karat, therefore 18 karat: 18/24 = .75 or 75% gold
:
let g = weight of the gold in the bracelet
let c = weight of copper
:
g + c = 15
c = (15-g)
and
19.3g + 9c = 238
replace c with (15-g)
19.3g + 9(15-g) = 238
19.3g + 135 - 9g = 238
19.3 - 9g = 238 - 135
10.3g = 103
g = 103/10.3
g = 10 grams of gold
and
15-10 = 5 grams of copper
:
Find the % gold; 10/15 = .667 or 67% gold, not 18k